OpenStudy (3psilon):

Why is 3^20 + 3^20 + 3^20 == 3^21 ??!? I'm so confused

OpenStudy (turingtest):

3^20+3^20+3^20 combine like terms... =3(3^20)=3^(20+1)=3^21

OpenStudy (anonymous):



OpenStudy (turingtest):

@3psilon let me know if you still don't get it...

OpenStudy (3psilon):

So you're factoring out the 3?

OpenStudy (3psilon):

Not sure how you go tto 3^(20+1)

OpenStudy (turingtest):

no, I am combining like terms call\[(3^{20})=x\]we then have\[(3^{20})+(3^{20})+(3^{20})=x+x+x=3x\]sub back in what we know x is...\[3x=3(3^{20})\]now we use the property\[x^a\cdot x^b=x^{a+b}\]and get\[3(3^{20})=3^1\cdot3^{20}=3^{20+1}=3^{21} \]

OpenStudy (3psilon):

OHHHH

OpenStudy (3psilon):

Yes! I thought I got it last night, but then I saw this problem and was embarrassed to ask again. Thank you turing

OpenStudy (turingtest):

very welcome! please never be too embarrassed to ask

OpenStudy (3psilon):

I won't! You will see me more often! haahaha have a good one